OpenStudy (anonymous):

This problem is Differentiation of Inverse Trigonometric Functions: y = arc sec squareroot(4x+1)

OpenStudy (anonymous):

\[y = arc \sec \sqrt{4x+1}\]

OpenStudy (turingtest):

\[\frac d{dx}\sec^{-1}u=\frac1{|u|\sqrt{1-u^2}}\frac {du}{dx}\]

OpenStudy (turingtest):

let\[u=\sqrt{4x+1}\]

OpenStudy (anonymous):

i need the answer so i will know if my answer is right.

OpenStudy (turingtest):

show me what you got so I can tell you

OpenStudy (anonymous):

i'm in the simplifying part already...